ADNAN MUHAMMED

Grade 12,

4 × 6 whole 1/3 is equal to 4 × 6 whole 1/3 is equal to

 4 ×  6 whole 1/3 is equal to                                   
 4 ×  6 whole 1/3 is equal to                                   

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

4 ×6 whole 1/3 is equal to 76/3

First we have to convert the mixed fraction into improper fraction 6 whole 1/3= 19/3 Then, 4 × 19/3

= 76/3


Thanks